OpenStudy (firejay5):

Is this question possible to solve?: While at Magic Springs, Katie goes down a frictionless water slide that is 9.0 m above the ground. At the bottom of the slide, Katie reaches a final speed of 13 m/s. What was Katie's initial speed when she pushed off the top of the slide?

OpenStudy (caozeyuan):

No, it can be done. use energy conservation to set up equations. pretty basic stuff

OpenStudy (firejay5):

Well you can't really solve it without mass

OpenStudy (caozeyuan):

but mass cancels right

OpenStudy (firejay5):

how

OpenStudy (caozeyuan):

\[\frac{ 1 }{ 2 }mv _{i} ^{2}+mgh=\frac{ 1 }{ 2 }mv _{f} ^{2}\]

OpenStudy (caozeyuan):

since m is non-zero, you can divide m on both sides

OpenStudy (caozeyuan):

there is really nothing tricky

OpenStudy (firejay5):

I got for \[v _{i}\] = 1.36m/s

OpenStudy (caozeyuan):

well, change against the official answer

OpenStudy (firejay5):

what do you mean

OpenStudy (caozeyuan):

I mean is the answer given to you? like at the back of the book or somewhere?

OpenStudy (firejay5):

not it's a worksheet and I solved it and I think it's right???

OpenStudy (firejay5):

@caozeyuan 1/2mvi^2 + mghi = 1/2mvf^2 + mghf 1/2mvi^2 + m(9.8)(9) = 1/2m(13^2) + m(9.8)(0) 1/2vi^2 + 88.2 = 84.5 + 0 0.5vi^2 = 3.7 vi^2 = 1.85 vi = 1.36 m/s

OpenStudy (caozeyuan):

this IS impossible to solve

OpenStudy (caozeyuan):

not because of the mass but becuase of the negative sign

OpenStudy (caozeyuan):

your vi is an imaginary number, therefore the speed don't exist

OpenStudy (firejay5):

I'll talk to my teacher about it, because he wrote it

OpenStudy (caozeyuan):

Yea, this is really werid

OpenStudy (firejay5):

I am sorry about the confusion, he's confusing sometimes